OpenStudy (anonymous):

A set of equations is given below: Equation C: y = 7x + 12 Equation D: y = 7x + 2 How many solutions are there to the given set of equations? One solution Two solutions Infinitely many solutions No solution

OpenStudy (anonymous):

@carolinar7

Vocaloid (vocaloid):

set the equations equal to each other 7x + 12 = 7x + 2 notice what happens when you subtract 7x from each side

OpenStudy (anonymous):

one solution

Vocaloid (vocaloid):

not quite 7x + 12 = 7x + 2 what do you get when you take away 7x from each side?

OpenStudy (anonymous):

no solution

Vocaloid (vocaloid):

can you explain how you got that answer?

OpenStudy (anonymous):

im guessing

Vocaloid (vocaloid):

7x + 12 = 7x + 2 subtract 7x from each side, please